### v1

# Raw Speech Enhancement with Deep State-Space Modeling

Abstract
We present aTENNuate, a simple deep state-space autoencoder configured for efficient online raw speech enhancement in an end-to-end fashion. The network’s performance is primarily evaluated on raw speech denoising, with additional assessments on tasks such as super-resolution and de-quantization. We benchmark aTENNuate on the VoiceBank + DEMAND and the Microsoft DNS1 synthetic test sets. The network outperforms previous real-time denoising models in terms of PESQ score, parameter count, MACs, and latency. Even as a raw waveform processing model, the model maintains high fidelity to the clean signal with minimal audible artifacts. In addition, the model remains performant even when the noisy input is compressed down to 4000Hz and 4 bits, suggesting general speech enhancement capabilities in low-resource environments.
Index Terms:
state-space models, autoencoder, denoising, super-resolution, de-quantization
I Introduction
Speech enhancement is crucial for improving both human-to-human communication, such as in hearing aids, and human-to-machine communication, as seen in automatic speech recognition (ASR) systems. The most challenging task of speech enhancement is removing background noises from speech signals. The complexity of speech patterns and the unknown characteristics of noise poses challenges, further exacerbated by the dense data samples in audio signals and the sensitivity of human perception to minor noise and distortions. [1] Traditional speech enhancement methods such as Wiener filtering [2], spectral subtraction [3] and principal component analysis [4] have shown satisfactory performance in stationary noise environments, but their effectiveness is often limited in non-stationary noise scenarios, resulting in artifacts such as musical noises and substantial degradation in both the quality and intelligibility of the enhanced speech [5].
Deep learning audio denoising methods are trained on large datasets of clean and noisy audio pairs, and attempt to capture the nonlinear relationship between the noisy and the clean signal features without prior knowledge of the noise statistics, as required by traditional denoising methods [6]. Speech features can be extracted from real or complex spectrograms of the noisy signal in the time-frequency domain, or the raw waveform [7].
Many state-of-the-art deep learning denoising models leverage the feature extraction capability of convolutional networks. The UNet convolutional encoder-decoder is a common network architecture in denoising models [8, 9, 10] exemplified by Deep Complex UNet [11] and the generative adversarial SEGAN [12]. However, CNNs lack the capabilities to model long-range temporal dependencies present in speech signals. Another class of models that are more capable in this regard are RNNs, examples of which are DCCRN [13] and FRCRN [14]. These discriminative models show limited robustness to different types of noise and generalization across diverse audio sources [15]. Likelihood-based generative models such as Denoising Diffusion Probabilistic Models (DDPMs) that treat denoising as a conditional generation problem have attempted to overcome the limitations of their discriminative counterparts in generalization to unseen situations [16]. DDPMs, along with Variational autoencoders [17] also belong to the unsupervised class of models promising enhanced generalization capability to unseen noise and acoustic conditions. These approaches, however, often result in a large number of parameters which are computationally expensive, and by their nonlinear recurrent nature cannot efficiently leverage parallel hardware (e.g. GPUs) for training.
Alternative approaches such as PercepNet [18] and RNNoise [19] have attempted to reduce network size and complexity by combining traditional speech enhancement methods with deep learning, resulting in smaller models with fewer parameters capable of running in real-time on general-purpose hardware. Similarly, methods that process raw waveform signals aim to maximize the expressive capabilities of deep networks without resorting to costly time-frequency conversions, as demonstrated by Facebook’s DEMUCS model’s real-time performance [20]. However, the majority of these models cannot perform real-time inference on general-purpose hardware, exemplified by Nvidia’s CleanUNet model [21]. Models such as DeepFilterNet [22] have leveraged speech-specific properties such as short-time speech correlations to achieve comparable results.
Here, we introduce the aTENNuate network, belonging to the class of Temporal Neural Networks (TENNs). It is a deep state-space model (SSM) [23, 24] optimized for real-time denoising of raw speech waveforms on the edge. By virtue of being a state-space model, the model is capable of capturing long-range temporal relationships present in speech signals, with stable linear recurrent units. Learning long-range correlations can be useful for capturing global speech patterns or noise profiles, and perhaps implicitly capture semantic contexts to aid speech enhancement performance [25].
During training of the aTENNuate network, we can use the infinite impulse response (IIR) kernels of the SSM layers as long convolutional kernels over the input features, which can be parallelized using techniques such as FFT convolution [26] or associative scan [27]. During inference, the temporal convolution layers can be converted into equivalent recurrent layers for efficient real-time processing on mobile devices, minimizing latency and reducing the need for excessive buffering of data. The evaluation code of a pre-trained aTENNuate network is available at our Github repo Brainchip-Inc/aTENNuate.
II State-space Modeling
In this section, we briefly describe what state-space models are, and recall how they can be configured for neural network processing, involving discretization and diagonalization. State-space models are general representations of linear time-invariant (LTI) systems, and they can be uniquely specified by four matrices: $A\in\mathbb{R}^{h\times h}$, $B\in\mathbb{R}^{h\times n}$, $C\in\mathbb{R}^{m\times h}$, and $D\in\mathbb{R}^{m\times n}$. The first-order ODE describing the LTI system is given as
$\dot{x}=Ax+Bu,\qquad y=Cx+Du,$ (1)
where $u\in\mathbb{R}^{n}$ is the input signal, $x\in\mathbb{R}^{h}$ is the internal state, and $y\in\mathbb{R}^{m}$ is the output. Here, we are letting $n>1,\,m>1$, which yields a multiple-input, multiple-output (MIMO) state-space model. For the remainder of this paper, we will ignore the $Du$ term as we do not use it.
The state-space model in its original form describes a continuous-time system, but in the field of digital signal processing, there are standard recipes for discretizing such a system into a discrete-time state-space model. One such method that we use in this work is the zero-order hold (ZOH), which gives us the discrete-time state-space matrices $\overline{A}$ and $\overline{B}$ as follows:
$\overline{A}=\exp(\Delta A),\qquad\overline{B}=(\Delta A)^{-1}\cdot(\exp(\Delta A)-1)\cdot\Delta B.$ (2)
The discrete state-space model is then given by
$x[t+1]=\overline{A}\,x[t]+\overline{B}\,u[t],\qquad y[t]=C\,x[t]$ (3)
In the context of recurrent neural networks (RNNs), this is essentially a linear RNN layer, which allows for efficient online inference and generation (in our case real-time speech enhancement), but at the same time efficient parallelization during training.
It is straightforward to check that the discrete-time impulse response is given as
$k[\tau]=C\,\overline{A}^{\tau}\,\overline{B},$ (4)
where $\tau$ denotes the kernel timestep. During training, $k$ can be considered the “full” long 1D convolutional kernel with shape (output channels, input channels, length), in the sense that the output $y$ can be computed via the long convolution $y_{j}=\sum_{i}u_{i}\ast k_{ij}$. By the convolution theorem, we can perform this operation in the frequency domain, which becomes a point-wise product $\hat{y}_{jf}=\sum_{i}u_{i}\hat{k}_{ijf}$. The hat symbol denotes the Fourier transform of the signal (with the index $f$ denoting the Fourier modes), which can be efficiently computed via Fast Fourier Transforms (FFTs).
It is a generic property (though not always true) that a diagonal form exists for the state-space model, meaning that we can almost always assume $\overline{A}$ to be diagonal [26], at the expense of potentially requiring $\overline{B}$ and $C$ to be complex matrices. Since the original system is a real system, the diagonal $\overline{A}$ matrix can only contain real elements and/or complex elements in conjugate pairs. (See Appendix A for an elementary demonstration of this fact using the impulse response.) In this work, we sacrifice a slight loss in expressivity by continuing to restrict $\overline{B}$ and $C$ to be real matrices and letting $A$ be a diagonal matrix with all complex elements (but not restricting them to come in conjugate pairs). Since we still want to work with real features [^1] This is a prior not required, as technically we can configure our network as complex-valued to handle complex features. However, we do not explore this configuration in this work., we then only take the real part of the impulse response kernel as such:
$k[\tau]=\Re(C\overline{A}^{\tau}\overline{B}),$ (5)
which equivalently in the state-space equation can be achieved by simply letting $y[t]=C\,\Re(x[t])$. This means that during online inference, we need to maintain the internal states $x$ as complex values, but only need to propagate their real parts to the next layer.
Similar to previous works in deep state-space modeling, we allow the parameters $\{A,B,C,\Delta\}$ to be directly learnable, which indirectly trains the kernel $k$. See Appendix C for more details. Unlike previous works in deep state-space modeling, we do not try to keep the sizes $n$, $h$, $m$ consistent, to allow for more flexibility in feature extraction at each layer, mirroring the flexibility in selecting channel and kernel sizes in convolutional neural networks [^2] The size of the internal state $h$, can be interpreted as the degree of parametrization of a basis temporal kernel, or some implicit (dilated) “kernel size” in the frequency domain. We explore this in a future work.. The flexibility of the tensor shapes requires us to carefully choose the optimal order of operations during training to minimize the computational load. This is a topic explored in previous work [28], and briefly discussed in Appendix C.
III Network Architecture
Figure 1: A schematic drawing of the network architecture, with only 2 encoder and decoder blocks shown for simplicity. The actual model has 6 encoder and decoder blocks. Note that there is no (spectral) processing on the input and output waveforms.
We use an hourglass network with long-range skip connections, similar in form to the Sashimi network [24] for audio generation. However, unlike previous works using state-space models for audio processing [24, 29, 30], our network directly takes in raw audio waveforms in the -1 to +1 range and outputs raw waveforms as well, with no one-hot encoding or spectral processing (e.g. STFT or iSTFT). Furthermore, we retain causality as much as possible for sake of real-time inference, meaning that we eschew any form of bidirectional state-space layers. See Fig. 1 for a schematic drawing.
As with typical auto-encoder networks, the audio features are down-sampled in the encoder and then up-sampled in the decoder. For the re-sampling operation, we use a simple operation that squeezes/expands the temporal dimension then projects the channel dimension [24]. More formally, for a reshaping ratio of $r$, a sequence of features can be down-sampled and up-sampled as:
[…]
The baseline network uses LayerNorm layers and SiLU activations. In addition, we include a “PreConv” layer which is a depthwise 1D convolution layer with a kernel size of 3, to enable better processing of local temporal features. The PreConv operation is omitted in the 2 SSM blocks in the neck and in any block with only one channel [^3] The neck operates on fully-downsampled features, and introducing PreConv layers will incur too much latency in real-time processing. We also omit PreConv layer in single-channel blocks, because it is too lossy.. See Appendix C for details on the computation of the theoretical latency for a given set of resampling factors and PreConv configurations. Using causal convolutions for PreConvs can eliminate additional latencies to the network [31], but we do not explore this implementation here. To better support mobile devices, we also test variants of the network with BatchNorm layers [^4] Since BatchNorm is a static form of normalization during inference, the normalization statistics and the affine parameters can be “folded” into the weights and biases of the previous layer, meaning that the layer does not need to be materialized during inference., ReLU activations, and omitting PreConv layers in the decoder (and omitting PreConv layers altogether). The results are reported in the next section.
IV Experiments
TABLE I: Comparing different variants of the aTENNuate network against other real-time audio denoising networks, in terms of performance, memory/computational requirements, and latency.
Model PESQ (Voicebank) PESQ (DNS1 no-reverb) Parameters MACs / sec Latency
DeepFilterNet3 [22] 3.16 2.58 2.13M 0.344G 40ms
DEMUCS [20] 2.56 2.65 33.53Ma 7.72Ga 40ms
PercepNet [18] 2.73b - 8.00M 0.80G 40ms
RNNoise [19] 2.43 1.94 0.06M 0.04G 20ms
aTENNuate (base) 3.27 2.98 0.84M 0.33G 46.5ms
     PreConvs only in encoder 3.21 2.84 0.84M 0.33G 31.25ms
     no PreConvs 3.06 2.59 0.84M 0.33G 16ms
       BatchNorm + ReLU 2.84 2.43 0.84M 0.33G 16ms
a These numbers are estimated by passing a one-second segment of data to the model.
b This metric is taken directly from the paper as an official implementation of the model does not exist.
For experiments, we train on the VCTK and LibroVox dataset, randomly mixed with noise samples from Audioset, Freesound, and DEMAND. We evaluate our denoising performance on the Voicebank + DEMAND testset, and the Microsoft DNS1 synthetic test set (with no reverberation) [32]. To guarantee no data leakage between the training and testing sets, we removed the clean and noise samples in the training set that were used to generate the synthetic testing samples. Both the input and output signals of the network are set at 16000 Hz. The loss function is a mix of SmoothL1Loss [33] and spectral loss at the ERB scale [22]. More details of the training pipeline are provided in Appendix B.
The evaluation metric is the average wideband PESQ score between the clean signals and the denoised outputs. The PESQ scores for different variants of the aTENNuate model against other real-time audio-denoising networks are reported in Table. I. We also report other network inference metrics including parameters, MACs, and latencies. For the latency metric, we focus on the theoretical latency of the network, not accounting for any processing latencies. In simpler terms, it is the maximum time range that the network needs to “wait” or look-forward to produce a denoised data point corresponding to the current input. As seen in the Table I, the PreConv layers (being depthwise) does not make much difference in terms of parameters and MACs, but does add considerably to latency. Unless otherwise denoted, we use the source code and pre-trained weights of each model and run it through a standardized PESQ evaluation pipeline.
For further inspection of the quality of the denoised samples produced by the network, we not only listen to the denoised samples (on both synthetic data and real recordings) but also provide a comparison of the denoised spectrogram and the clean spectrogram. This is to ensure that the denoised samples do not contain any unnatural artifacts that are common with raw audio processing systems [^5] For example, the network may generate low-frequency artificial signals, resulting in a background “humming” noise. It may also attempt to aggressively attenuate high frequency speech components, resulting in a robotic tone of speech., which may not be captured by the PESQ score [34]. See Fig. 2 for the spectrograms.
Figure 2: A comparison of the spectrograms among the noisy signal, the clean ground truth, and the denoised output of a sample in the DNS1 synthetic testset (no reverb). Besides a minor low-frequency artifact in the silent region, the denoised output matches very close to the ground truth signal, despite not using any pre/post-processing in the spectral domain.
In addition to audio denoising, we also perform studies on the ability of our network to perform super-resolution and de-quantization on highly compressed data. This involves intentionally performing down-sampling and quantization of the input signals, in that order, and re-training the network to handle the degraded/compressed inputs [^6] It is possible to perform parameter-efficient fine-tuning of a baseline model instead, such as using low-rank adaptations on the state-space matrices. We leave this to future work.. To use the same network architecture to interface with down-sampled audios, we perform interleaved repeats of the input signals to restore the original sample rate [^7] An alternative is to simply remove an encoder block with the same downsampling factor as the downsampled input.. For quantization, we perform mu-law encoding on the input signals down to the desired bitwidth, then rescale the quantized signal back to the -1 to +1 range. The super-resolution and de-quantization results are reported in Table. II. Note that the outputs are still evaluated against clean signals at 16000 Hz and full precision.
TABLE II: The average PESQ scores of the model outputs when the noisy inputs are down-sampled and quantized.
Input Type VoiceBank DNS1
8000 Hz & 8 bit 3.19 2.88
4000 Hz & 8 bit 3.04 2.72
8000 Hz & 4 bit 2.90 2.55
4000 Hz & 4 bit 2.72 2.39
V Future Directions
To make the network even more mobile-friendly, we plan to explore the sparsification and quantization of the network weights and activations. Potentially, we can study a low-rank realization of the state-space matrices, which form the majority of the parameter count. State-space models are also known to be easily adaptable for a spiking implementation on neuromorphic hardware, so it will also be interesting to see whether our model admits an efficient spiking neural network realization, which will further reduce the power required for the solution.
VI Conclusion
We introduced a lightweight deep state-space autoencoder, aTENNuate, that can perform raw audio denoising, super-resolution, and de-quantization. Compared to previous works, the key features of this network are: 1) consisting of state-space layers that can be efficiently trained and configured for inference, 2) allowing for real-time inference with low latency, 3) architecturally simple and light in parameters and MACs, 4) capable of processing raw audio waveforms directly without requiring pre/post-processing, and 5) highly competitive with other speech enhancement solutions.
VII Acknolwedgement
We thank Temi Mohandespour and Keith Johnson for contributing to the early stages of the project. We also thank M. Anthony Lewis, Douglas McLelland, Kristofor Carlson, and Chris Jones for providing useful feedback on the manuscript.
Appendix A Diagonalization of the $\overline{A}$ matrix
If we allow $\overline{B}$ and $C$ to be complex matrices, then we can assume $\overline{A}$ to be diagonal without any loss of generality. To see why, we simply let $\overline{A}=P^{-1}\Lambda_{A}P$ (where $\Lambda_{A}$ is the diagonalized $\overline{A}$ matrix, and $P$ is the similarity matrix) and observe the following:
$\begin{split}\forall t,\qquad k[t]&=C(P^{-1}\Lambda_{A}P)^{t-1}\overline{B}\\ &=CP^{-1}\underbrace{\Lambda_{A}(PP^{-1})\,...\,\Lambda_{A}(PP^{-1})}_{\text{repeat $t-1$ times}}P\overline{B}\\ &=(CP^{-1})\Lambda_{A}^{t-1}(P\overline{B})\\ &=C^{\prime}\Lambda_{A}^{t-1}\overline{B}^{\prime},\end{split}$ (7)
where $\overline{B}^{\prime}$ and $C^{\prime}$ are complex matrices that have “absorbed” the similarity matrix $P$, but WLOG we can just redefine them to be $\overline{B}$ and $C$. Since $\overline{A}$ is a real matrix, the complex eigenvalues in $\Lambda_{A}$ must come in conjugate pairs. And WLOG we can again redefine $\Lambda_{A}$ as $\overline{A}$.
Appendix B Experiment Details
Our baseline model is trained with:
  • 500 epochs
  • AdamW optimizer with the PyTorch default configs
  • A cosine decay scheduler with a linear warmup period equal to 0.01 of the total training steps, updating after every optimizer step
  • gradient clip value of 1
  • layer normalization (over the feature dimension) with elementwise affine parameters
  • SiLU activation
  • no dropout
The high-level training pipeline for the raw audio denoising model is to simply generate synthetic noisy audios by randomly mixing clean and noise audio sources. The noisy sample is then used as input to the model, and the clean sample is used as the target output.
For the clean training samples, we use the processed VCTK and LibriVox datasets that can be downloaded from the Microsoft DNS4 challenge. We also use the noise training samples from the DNS4 challenge as well, which contains the Audioset, Freesound, and DEMAND datasets. For all audio samples, we use the librosa library to resample them to 16 kHz and load them as numpy arrays.
For the LibriVox audio samples which form long continuous segments of human subjects reading from a book, we simply concatenate all the numpy arrays, and pad at the very end such that the array can be reshaped into (segments, $(\text{segments},\,2^{17})$. For all the other audio samples consisting of short disjoint segments, we perform intermediate paddings when necessary, to ensure a single recording does not span two rows in the final array. For audio samples longer than length $2^{17}$, we simply discard them. The input length to our network during training is then also $2^{17}$.
For every epoch, we use the entirety of the VCTK dataset and 10 percent of a randomly sampled subset of the LibriVox dataset. For each clean segment, we pair it with a randomly sampled noise segment (with replacement). The clean and noise samples are added together with an SNR sampled from -5 dB to 15 dB, and the synthesized noisy sample is then rescaled to a random level from -35 dB to -15 dB. Furthermore, we perform random temporal and frequency masking (part of the SpecAugment transform) on only the input noisy samples.
For the loss function, we combine SmoothL1Loss with spectral loss on the ERB scale, without any equalization procedure on the raw waveforms or spectrograms. The $\beta$ parameter of the SmoothL1Loss is set at 0.5, and the spectral loss is weighted by a factor that grows from 0 to 1 linearly during training. We use a learning rate of 0.005 and a weight decay of 0.02.
Appendix C Network Details
C-A Initialization of SSM Parameters
As mentioned in Section II, we make the SSM parameters $\{A,B,C,\Delta\}$ trainable. Recall that we take $A\in\mathbb{C}^{h}$ to be a complex vector (or a complex diagonal matrix). For stability, we treat the real and complex parts of $A$ separately. The real part $\Re(A)$ is parameterized as $-\text{softplus}(a_{r})$, where $a_{r}$ is initialized with the value $-0.4328$, giving $\Re(A)=-1/2$ initially. Note that due to the positivity of softplus, $\Re(A)$ will always remain negative during training, which ensures the stability of the SSM layer. The imaginary part $\Im(a)$ is parameterized directly and initialized with $\pi(i-1)$ where $i$ is the state index. The matrix $B\in\mathbb{R}^{h\times n}$ is initialized with all ones, and the matrix $C\in\mathbb{R}^{m\times h}$ is initialized with Kaiming normal random variables (assuming a fan in of $h$). Finally, we initialize $\Delta$ with $0.001\times 100^{\lfloor i/16\rfloor}$, giving a series of geometrically spaced values from 0.001 to 0.1 in blocks of 16. These initializations and parameterizations are not absolutely required, and we suspect that any reasonable approach respecting the stability of the SSM layer will suffice.
C-B Optimal Contraction Order
If we let $K(t)=\overline{A}^{t}$ be the “basis kernels” of the SSM layer, and its Fourier transform be $\hat{K}(f)$, then in einsum form, the SSM layer operations during training can be expressed as
$\hat{y}_{bjf}=\hat{x}_{bif}\overline{B}_{ni}\hat{K}_{nf}C_{jn},$ (8)
where $\{b,i,j,n,f\}$ indexes the batch size, input channels, output channels, internal states, and Fourier modes respectively. With abuse of notation, we similarly let $\{B,I,J,N,F\}$ be the sizes of the five aforementioned dimensions. Note that the number of Fourier modes $F$ is the same as the length of the signal $L$.
There are two main ways to compute $\hat{y}$. First, we can perform the operations of Eq. 8 from left to right normally, corresponding to projecting the input, performing the FFT convolution, then projecting the output. Alternatively, we can compute the full kernel, then perform the full FFT convolution with the input as $\hat{x}_{bif}(\overline{B}_{ni}\hat{K}_{nf}C_{jn})$. If we only focus on the computational requirements of the forward pass, then the first contraction order will result in $BNIF+BNF+BJNF\approx BNF(I+J)$ units of computation. The second contraction order will result in $JNI+JNIF+BJIF\approx JIF(B+N)$ units of computation. Therefore, we see that the optimal contraction order is intimately linked with the dimensions of the tensor operands. More formally, the first order of contraction is more optimal only when $BNF(I+J)<JIF(B+N)$ or $\frac{1}{B}+\frac{1}{N}>\frac{1}{I}+\frac{1}{J}$.
C-C Network Latency
As mentioned in Section III and Fig. 1 of the main text, our model is an autoencoder network with long-range skip connections between the blocks of encoder and decoder. Each SSM layer retains the length and the channels of the input features, and a resampling layer performs both the temporal resampling and channel projection operations. For example, the first SSM block maps the input features as $(1,L)\to(1,L)$, followed by a resampling layer that maps the features as $(1,L)\to(16,L/4)$, through the process described in Section III. The resampling factor and output channel of each block are reported in Table III. For all the blocks, the hidden state of the SSM layer is fixed to $h=256$.
If the network does not have any (non-causal) convolutional layers (or PreConvs), then the theoretical latency can be simply determined from the product of the resampling factors, since the “stride sizes” is the same as the “kernel sizes” for the these layers. In the case of the network given in Table III, the product of the resampling factors (or the total resampling factor) is 256, and the sample rate of the input signal is 16000Hz, so the latency of the network is at least $256/16000\text{Hz}=16$ milliseconds [^8] Technically, it should be $(256-1)/16000\text{Hz}=15.9\text{ms}$, but this makes little difference.. In the presence of PreConvs, or centered convolutional layers of kernel sizes of 3, then there are additional latencies introduced due to the non-causal nature of the convolutions. More speficially, the additional latency is the look-forward timestep of the kernel, or simply the step size of the input features.
TABLE III: Resampling factor and output channel of each block of the network, which consists of an encoder performing down-samplings, an intermediate bottleneck, a decoder performing up-samplings, and finally an output processor.
Layers Resample Factor Channels PreConv latency
Encoder
   Block 1 4 16
   Block 2 4 32 0.25ms
   Block 3 2 64 1ms
   Block 4 2 96 2ms
   Block 5 2 128 4ms
   Block 6 2 256 8ms
Neck
   Block 1 1 256
   Block 2 1 256
Decoder
   Block 1 2 128 8ms
   Block 2 2 96 4ms
   Block 3 2 64 2ms
   Block 4 2 32 1ms
   Block 5 4 16 0.25ms
   Block 6 4 1
Output
   Block 1 1 1
   Block 2 1 1